# Zero modes and index theorems for non-Hermitian Dirac fermions

### v1

@@ III Jackiw-Rebbi index theorem
The Jackiw-Rebbi index theorem applies to one-dimensional massive Dirac fermions for which the mass order $m(x)$, accompanied by a Hermitian matrix that fully anticommutes with the free Dirac Hamiltonian assumes the profile of a domain-wall, given by $m(x\to\pm\infty)=\pm m_{0}$ (constant) otherwise arbitrary [7]. In one-dimensional NH Dirac system the corresponding NH Dirac operator can be obtained by taking $\Gamma_{1}=\sigma_{1}$ and $M=\sigma_{3}$, and the mass order is represented by the Hermitian operator $m(x)\sigma_{3}$. The corresponding total massive NH Dirac operator reads as
@@
Before I delve into the search for the zero modes in the spectrum of the above NH operator, it is worth pausing to discuss the choice of the Hermitian matrix for the mass order. For this purpose and without any loss of generality, I ignore any spatial modulation of the mass order and assume it to be constant $m(x)=m_{0}$ for any $x$. The one-dimensional Dirac Hamiltonian $\sigma_{1}k_{x}$ supports two mass matrices, namely $\sigma_{2}$ and $\sigma_{3}$. Once I choose $M=\sigma_{3}$ in Eq. (2) to construct the NH Dirac operator, only for one mass matrix $\sigma_{3}$ one can find a real definite eigenvalue spectrum for $|\alpha|<1$ of the corresponding total NH massive Dirac operator given by $(\sigma_{1}+i\alpha\sigma_{2})k_{x}+\sigma_{3}m_{0}$. The eigenvalues of this NH operator are given by $\pm\sqrt{(1-\alpha^{2})k^{2}_{x}+m^{2}_{0}}$. Such mass order is called ‘commuting class mass’ as $M,\sigma_{3}]=0 [$35]. In this section and the next one, I only focus on such ‘commuting class mass’ orders.
Now I return to finding the zero-energy bound state in the spectrum of the NH operator shown in Eq. (17). In the Hermitian system, the corresponding operator satisfies a unitary particle-hole symmetry, generated by $\sigma_{2}$ as $\{H^{\rm NH}_{\rm Dir}(\bm{k}\to-i{\bm{\nabla}},0),\sigma_{2}\}=0$. Hence, any mode at precise zero-energy is an eigenstate of $\sigma_{2}$. However, in the NH setup $\sigma_{2}$ generates the pseudo particle-hole symmetry [41, 42], given by
@@
The pseudo particle-hole symmetry guarantees that if there exists a right eigenvector for the zero-energy mode then there must also exists its dual the left eigenvector. With this symmetry in hand I proceed to find the right zero-energy eigenvector $\left|\Psi_{0}\right>_{R}=(u_{R},v_{R})^{\top}$ in the spectrum of NH massive Dirac fermions with a domain-wall mass by solving the following differential equation
@@
Notice that the bound state in the presence of a domain-wall for $|\alpha|<1$, becomes a plane-wave solution at zero energy for $|\alpha|>1$. The existence of such a zero-energy plane-wave solution in the presence of an underlying domain-wall is only possible in NH Dirac systems. Next, I proceed to scrutinize the Jackiw-Rossi index theorem in two dimensions in NH Dirac systems.
@@ IV Jackiw-Rossi index theorem
The Jackiw-Rossi index theorem dictates the number of zero-energy bound states in the spectrum of two-dimensional massive Dirac fermions, when two mutually anticommuting Dirac masses with respective amplitudes of $\Delta_{1}$ and $\Delta_{2}$ and a requisite U(1) symmetry between them feature vortex-like topological defects in real space. Here, I extend this index theorem for NH Dirac fermions for which the corresponding operator is constructed following the general principle from Eq. (2). Before delving into such real space topological defects for the composite (two-component) mass order, consider the corresponding NH operator in the presence of two uniform masses that will allow us to uniquely identify the mass matrix $M$, appearing in the anti-Hermitian component of the NH Dirac operator in Eq. (2), given by
@@
The Dirac Hamiltonian for free fermions in two dimensions involves only two mutually anticommuting $\Gamma$ matrices, namely $\Gamma_{1}$ and $\Gamma_{2}$, and the theory enjoys an SU(2) chiral symmetry, generated by $\{\Gamma_{34},\Gamma_{45},\Gamma_{53}\}$, where $\Gamma_{jk}=i\Gamma_{j}\Gamma_{k}$ [44, 45]. The two-dimensional four-component Dirac system altogether supports four mass orders. Three of them break the SU(2) chiral symmetry and constitute the set of chiral symmetry breaking mass orders, explicitly given by $\{\Gamma_{3},\Gamma_{4},\Gamma_{5}\}$. Specifically in two dimensions there exists a fourth mass order, represented by $\Gamma_{12}$, which transforms as a scalar under the chiral rotation. I name it a ‘chiral scalar’ mass, which, however, breaks the time-reversal symmetry that in the announced representation in Eq. (IV) is generated by the anti-unitary operator ${\mathcal{T}}=\Gamma_{14}{\mathcal{K}}$, where $\mathcal{K}$} is the complex conjugation. The chiral symmetry breaking mass orders preserve the time-reversal symmetry. While all the four mass orders anticommute with the free Dirac Hamiltonian, the chiral scalar mass operator commutes with all the members of the chiral symmetry breaking mass order. This observation leads to a unique choice of $M=\Gamma_{12}$ in Eq. (22) in the construction of the NH Dirac operator such that two mass order matrices $\Gamma_{3}$ and $\Gamma_{4}$ belong to the commuting class mass family for NH Dirac fermions and the eigenspectrum of the total NH massive Dirac operator is given by $\pm E_{\alpha}(\bm{k})$, where
@@
Although not directly relevant at this stage, with a suitable definition of the spinor basis appropriate for monolayer graphene, a prototypical example of Dirac systems, the chiral symmetry breaking mass orders corresponds to the charge-density-wave ($\Gamma_{5}$[46] and two, namely the real ($\Gamma_{3}$) and imaginary ($\Gamma_{4}$), components of the Kekulé valence bond order [45, 26], while the time-reversal symmetry breaking chiral scalar mass corresponds to the Haldane’s quantum anomalous Hall insulator [47]. Furthermore, the same NH Dirac Hamiltonian in an appropriate Nambu-doubled spinor basis also describes the gapless surface states of three-dimensional NH topological insulators, devoid of any NH skin effect [48]. In that case, $M=\Gamma_{12}$ corresponds to the surface magnetization in the $z$ direction, and $\Delta_{1}$ ($\Delta_{2}$) is the real (imaginary) component of an $s$-wave pairing [25]. However, the following outcomes are insensitive to these details.
@@
Finding the solution of the above equation becomes much more efficient by noticing the fact that $\{H^{\rm NH}_{\rm Dir}(\bm{k}\to-i{\bm{\nabla}},\alpha,n),\Gamma_{5}\}=0$. Hence, any mode that is bound at zero energy must then be an eigenstate of $\Gamma_{5}={\rm diag}.(1,-1,-1,1)$. Furthermore, I note that in the Hermitian limit ($\alpha=0$), $\{H^{\rm NH}_{\rm Dir}(\bm{k}\to-i{\bm{\nabla}},0,n),{\mathcal{A}}\}=0$ where ${\mathcal{A}}=\Gamma_{23}{\mathcal{K}}$ is the generator of the anti-unitary particle-hole symmetry [49]. Therefore, the zero-energy modes are also eigenstates of $\mathcal{A}$} in the Hermitian system. On the other hand, for any non-trivial $\alpha$ I find
@@
Therefore, the generator of the anti-unitary particle-hole symmetry of Hermitian Jackiw-Rossi model connects the left and right zero-energy eigenmodes (when exist) of its NH incarnation, yielding its pseudo particle-hole symmetry. With these spectral symmetries of the Hermitian and NH Jackiw-Rossi operators in hand, I now proceed to find the explicit solutions for the zero-energy modes.
In what follows, I consider a vortex-type defect, characterized by an integer $n>0$. The solutions for an anti-vortex with $n<0$ can readily be obtained from the solutions below by noting that the operator (Hermitian or NH) for the anti-vortex is obtained after a unitary rotation of the operator for vortex by $\Gamma_{45}$. Hence, zero mode solutions for anti-vortex are obtained after unitarily rotating the ones for a vortex-type defect by $\Gamma_{45}$.
The zero-energy right eigenmodes localized near the core of the vortex whose amplitude falls off as $r\to\infty$ and square integrable with respect to its bi-orthogonal product with the left eigenmode near the origin as $r\to 0$ is then found to be of the following generic form
$\left|\Psi_{0}(r,\varphi)\right>_{R}=C\left(\begin{array}[]{c}0\\ v^{R}_{1}(r,\varphi)\\ u^{R}_{2}(r,\varphi)\\ 0\end{array}\right).$ (32)
@@
where $c_{0}^{f}$, $c_{0}^{g}$, and $c_{\infty}$ are three arbitrary constants that can be determined by matching the solutions at an intermediate value of $r$. For square normalizability of the solutions near the origin the integer values of $m$ (required for the single-valued solutions) are bounded within the range $1/2\leq m\leq n-1/2$, yielding $m=0,1,\cdots,n-1$. Therefore, there are a total of $n$ number of allowed values of $m$, leading to a total of $n$ number of zero-energy solutions, which is the celebrated Jackiw-Rossi index theorem [10, 11]. Also notice that for the zero mode solutions to be normalizable as $r\to\infty$, the NH parameter must satisfy $|\alpha|<1$. Therefore, a NH planar massive Dirac system continues to foster $n$ number of normalizable zero-energy modes in the presence of a U(1) mass vortex, as long as the effective Fermi velocity of the collection of NH Dirac fermions remain purely real. When $|\alpha|>1$, the bound states I found otherwise for $|\alpha|<1$, become plane-waves at zero energy. Notice that existence of such zero-energy plane-wave solutions in the presence of an underlying vortex-type defect solely results from the non-Hermiticity in the system.
When $n$ is an odd integer, there exists one special zero mode solution with m=(n1)/2m=(n-1)/2, for which a closed solution can be found for any arbitrary $r$ and the corresponding right eigenvector given by
@@
This special zero-energy mode will play an important role when I search for the zero modes in the spectrum of a generalized NH Jackiw-Rossi operator, constructed by exploiting its pseudo particle-hole symmetry, see Eq. (27).
@@ IV.1 Generalized NH Jackiw-Rossi operator
The spectral symmetry of the Jackiw-Rossi Hamiltonian with respect to the antiunitary operator ${\mathcal{A}}=\Gamma_{23}{\mathcal{K}}$, allows one to introduce additional terms therein, leading to the generalized Jackiw-Rossi Hamiltonian. When extended to the case of NH systems, the generalized NH Jackiw-Rossi operator takes the following form
@@
Notice that that the matrix operator $\Gamma_{34}$, appearing with the minimally coupled gauge field $\bm{A}$, causes the U(1) rotation between the real and imaginary components of the composite mass order, represented by $\Gamma_{3}$ and $\Gamma_{4}$, respectively. Hence, the presence of such gauge field is necessary for the stability of an isolated or deconfined vortex. The physical origin of such a gauge field and two other quantities, namely $\mu$ and $h$, depends on microscopic details. For example, in graphene-based Dirac system of spinless fermions the gauge field enters as the axial or chiral gauge field pointing in the opposite directions near two complementary valleys, and $\mu$ ($h$) corresponds to Haldane’s quantum anomalous Hall insulator mass (chiral chemical potential). On the other hand, on the surface of three-dimensional topological insulators featuring a superconducting vortex, the gauge field results from regular electromagnetic vector potential, and $\mu$ ($h$) represents the $z$-component of surface magnetization (chemical potential). The forthcoming discussion in this section on the existence of zero-energy modes in the spectrum of $H^{\rm Gen}_{\rm JR}$ will be guided by the existing results in the Hermitian system ($\alpha=0$[50].
@@
The dependence of the zero modes on $\mu$ and $h$ are captured by the functions $f(r)$ and $g(r)$, which are respectively given by
@@
When $\mu>h$, as the Bessel functions are well-behaved near the origin and as $r\to\infty$, the normalization of the zero mode solutions are solely determined by the exponential factors in Eq. (IV.1), yielding normalizable solutions only if $|\alpha|<1$, i.e., when the effective Fermi velocity of the collection of NH Dirac fermions is real. Notice that when $h>\mu$, as the modified Bessel functions grow exponentially as $r\to\infty$, one can find normalizable solutions for zero modes if and only if $\Delta^{2}_{0}+\mu^{2}\geq h^{2}$ and when $|\alpha|<1$. The former condition is same as in Hermitian systems. Finally, I note that for $|\alpha|>1$, $J_{n}\leftrightarrow I_{n}$ and one cannot find any normalizable solution irrespective of the relative strength of $\mu$ and $h$.
@@
On the other hand, inside the vortex core the functions $f(r)$ and $g(r)$ take the form shown in Eqs. (49) and (50), respectively, with $n=1$ and $\mu^{2}-1/(2\xi^{2})\to\mu^{2}$ therein. Even in this case, I find that normalizable zero energy modes exists only when $|\alpha|<1$, i.e., the Fermi velocity of the collection of NH Dirac fermions is purely real.

### v2

@@ III Jackiw-Rebbi index theorem
The Jackiw-Rebbi index theorem applies to one-dimensional massive Dirac fermions for which the mass order $m(x)$, accompanied by a Hermitian matrix that fully anticommutes with the free Dirac Hamiltonian assumes the profile of a domain wall, given by $m\;(x\to\pm\infty)=\pm m_{0}$ (constant) otherwise arbitrary [7]. In one-dimensional NH Dirac systems, the corresponding NH Dirac operator can be obtained by taking $\Gamma_{1}=\sigma_{1}$ and $M=\sigma_{3}$, and the mass order is represented by the Hermitian operator $m(x)\sigma_{3}$. The corresponding total massive NH Dirac operator then reads as
@@
Before I delve into the search for the zero modes in the spectrum of the above NH operator, it is worth pausing to discuss the choice of the Hermitian matrix for the mass order. For this purpose and without any loss of generality, I ignore any spatial modulation of the mass order and assume it to be constant or uniform $m(x)=m_{0}$ for any $x$. The one-dimensional Dirac Hamiltonian $\sigma_{1}k_{x}$ supports two mass matrices, namely $\sigma_{2}$ and $\sigma_{3}$. Once I choose $M=\sigma_{3}$ in Eq. (2) to construct the NH Dirac operator, only for one mass matrix $\sigma_{3}$ one can find a real definite eigenvalue spectrum for $|\alpha|<1$ of the corresponding total NH massive Dirac operator given by $(\sigma_{1}+i\alpha\sigma_{2})k_{x}+\sigma_{3}m_{0}$. Explicitly, the eigenvalues of this NH operator are given by $\pm\sqrt{(1-\alpha^{2})k^{2}_{x}+m^{2}_{0}}$. Such mass order is called “commuting class mass” as $M,\sigma_{3}]=0 [$38]. In this Section and the next one, I only focus on such “commuting class mass” orders.
Now I return to finding the zero-energy bound state in the spectrum of the NH operator shown in Eq. (17). In the Hermitian system, the corresponding operator satisfies a unitary particle-hole symmetry, generated by $\sigma_{2}$ as $\{H^{\rm NH}_{\rm Dir}(\bm{k}\to-i{\bm{\nabla}},0),\sigma_{2}\}=0$. Hence, any mode at precise zero energy is an eigenstate of $\sigma_{2}$. However, in the NH setup $\sigma_{2}$ generates the pseudo-particle-hole symmetry [44, 45], given by
@@
The pseudo-particle-hole symmetry guarantees that if there exists a right eigenvector for the zero-energy mode then there must also exist its dual the left eigenvector. With this symmetry in hand I proceed to find the right zero-energy eigenvector $\left|\Psi_{0}\right>_{R}=(u_{R},v_{R})^{\top}$ in the spectrum of NH massive Dirac fermions with a domain-wall mass by solving the following differential equation
@@
Notice that the bound state in the presence of a domain wall for $|\alpha|<1$ becomes a plane-wave solution at zero energy for $|\alpha|>1$. The existence of such a zero-energy plane-wave solution in the presence of an underlying domain wall is only possible in NH Dirac systems. Next, I proceed to scrutinize the Jackiw-Rossi index theorem in two dimensions in NH Dirac systems.
@@ IV Jackiw-Rossi index theorem
The Jackiw-Rossi index theorem dictates the number of zero-energy bound states in the spectrum of two-dimensional massive Dirac fermions, when two mutually anticommuting Dirac masses with respective amplitudes of $\Delta_{1}$ and $\Delta_{2}$ and a requisite U(1) symmetry between them feature vortexlike topological defects in real space. Here, I extend this index theorem for NH Dirac fermions for which the corresponding operator is constructed following the general principle from Eq. (2). Before delving into such real space topological defects for the composite (two-component) mass order, consider the corresponding NH operator in the presence of two uniform masses that will allow us to uniquely identify the mass matrix $M$, appearing in the anti-Hermitian component of the NH Dirac operator in Eq. (2), given by
@@
The Dirac Hamiltonian for free fermions in two dimensions involves only two mutually anticommuting $\Gamma$ matrices, namely $\Gamma_{1}$ and $\Gamma_{2}$, and the theory enjoys an SU(2) chiral symmetry, generated by $\{\Gamma_{34},\Gamma_{45},\Gamma_{53}\}$, where $\Gamma_{jk}=i\Gamma_{j}\Gamma_{k}$ [48, 49]. The two-dimensional four-component Dirac system altogether supports four mass orders. Three of them break the SU(2) chiral symmetry and constitute the set of chiral symmetry breaking mass orders, explicitly given by $\{\Gamma_{3},\Gamma_{4},\Gamma_{5}\}$. Specifically in two dimensions there exists a fourth mass order, represented by $\Gamma_{12}$, which transforms as a scalar under the chiral rotation. I name it a “chiral scalar” mass, which, however, breaks the time-reversal symmetry that in the announced representation in Eq. (IV) is generated by the antiunitary operator ${\mathcal{T}}=\Gamma_{14}{\mathcal{K}}$, where $\mathcal{K}$} is the complex conjugation. The chiral symmetry breaking mass orders preserve the time-reversal symmetry. While all the four mass orders anticommute with the free Dirac Hamiltonian, the chiral scalar mass operator commutes with all the members of the chiral symmetry breaking mass order. This observation leads to a unique choice of $M=\Gamma_{12}$ in Eq. (22) in the construction of the NH Dirac operator such that two mass order matrices $\Gamma_{3}$ and $\Gamma_{4}$ belong to the “commuting class mass” family for NH Dirac fermions. The eigenspectrum of the total NH massive Dirac operator is the given by $\pm E_{\alpha}(\bm{k})$, where
@@
Although not directly relevant at this stage, it is worth noting that with a suitable definition of the spinor basis appropriate for monolayer graphene, a prototypical example of Dirac systems, the chiral symmetry breaking mass orders correspond to the charge-density-wave ($\Gamma_{5}$[50] and two, namely the real ($\Gamma_{3}$) and imaginary ($\Gamma_{4}$), components of the Kekulé valence bond order [49, 27], while the time-reversal symmetry breaking chiral scalar mass corresponds to the Haldane’s quantum anomalous Hall insulator [51]. Furthermore, the same NH Dirac Hamiltonian in an appropriate Nambu-doubled spinor basis also describes the gapless surface states of three-dimensional NH topological insulators, devoid of any NH skin effect [52]. In that case, $M=\Gamma_{12}$ corresponds to the surface magnetization in the $z$ direction, and $\Delta_{1}$ ($\Delta_{2}$) is the real (imaginary) component of an $s$-wave pairing [26]. However, the following outcomes are insensitive to these details.
@@
Finding the solution of the above equation becomes much more efficient by noticing the fact that $\{H^{\rm NH}_{\rm Dir}(\bm{k}\to-i{\bm{\nabla}},\alpha,n),\Gamma_{5}\}=0$. Hence, any mode that is bound at zero energy must then be an eigenstate of $\Gamma_{5}={\rm diag}.(1,-1,-1,1)$. Furthermore, I note that in the Hermitian limit ($\alpha=0$), $\{H^{\rm NH}_{\rm Dir}(\bm{k}\to-i{\bm{\nabla}},0,n),{\mathcal{A}}\}=0$ where ${\mathcal{A}}=\Gamma_{23}{\mathcal{K}}$ is the generator of the antiunitary particle-hole symmetry [53]. Therefore, the zero-energy modes are also eigenstates of $\mathcal{A}$} in the Hermitian system. On the other hand, for any nontrivial $\alpha$, I find
@@
Therefore, the generator of the antiunitary particle-hole symmetry of Hermitian Jackiw-Rossi model connects the left and right zero-energy eigenmodes (when they exist) of its NH incarnation, yielding its pseudo-particle-hole symmetry. With these spectral symmetries of the Hermitian and NH Jackiw-Rossi operators in hand, I now proceed to find the explicit solutions for the zero-energy modes.
In what follows, I consider a vortex-type defect, characterized by an integer $n>0$. The solutions for an antivortex with $n<0$ can readily be obtained from the solutions below by noting that the operator (Hermitian or NH) for the antivortex is obtained after a unitary rotation of the operator with a vortex-type defect by $\Gamma_{45}$. Hence, zero mode solutions for the antivortex are also obtained after unitarily rotating the ones for a vortex-type defect by $\Gamma_{45}$, which I discuss next.
The zero-energy right eigenmodes localized near the core of the vortex whose amplitude falls off as $r\to\infty$ and square integrable with respect to its biorthogonal product with the left eigenmode near the origin as $r\to 0$ is then found to be of the following generic form
$~\left|\Psi_{0}(r,\varphi)\right>_{R}=C\left(\begin{array}[]{c}0\\ v^{R}_{1}(r,\varphi)\\ u^{R}_{2}(r,\varphi)\\ 0\end{array}\right).$ (32)
@@
where $c_{0}^{f}$, $c_{0}^{g}$, and $c_{\infty}$ are three arbitrary constants that can be determined by matching the solutions at an intermediate value of $r$. For square normalizability of the solutions near the origin the integer values of $m$ (required for the single-valued solutions) are bounded within the range $1/2\leq m\leq n-1/2$, yielding $m=0,1,\cdots,n-1$. Therefore, there are a total of $n$ number of allowed values of $m$, leading to a total of $n$ number of zero-energy solutions, which is the celebrated Jackiw-Rossi index theorem, originally announced in Hermitian systems [10, 11]. Also notice that for the zero mode solutions to be normalizable as $r\to\infty$, the NH parameter must satisfy $|\alpha|<1$. Therefore, a NH planar massive Dirac system continues to foster $n$ number of normalizable zero-energy modes in the presence of a U(1) mass vortex, as long as the effective Fermi velocity of the collection of NH Dirac fermions remains purely real. When $|\alpha|>1$, the bound states I found otherwise for $|\alpha|<1$, become plane waves at zero energy. Existence of such zero-energy plane-wave modes in the presence of an underlying vortex-type defect solely results from the non-Hermiticity in the system.
When $n$ is an odd integer, there exists one special zero mode solution with m=(n1)/2m=(n-1)/2, for which a closed solution can be found for any arbitrary $r$. The corresponding right eigenvector is given by
@@
This special zero-energy mode will play an important role when I search for the zero modes in the spectrum of a generalized NH Jackiw-Rossi operator, constructed by exploiting its pseudo-particle-hole symmetry, see Eq. (27), which I discuss next.
@@ IV.1 Generalized NH Jackiw-Rossi operator
The spectral symmetry of the Jackiw-Rossi Hamiltonian with respect to the antiunitary operator ${\mathcal{A}}=\Gamma_{23}{\mathcal{K}}$, allows one to introduce additional terms therein, leading to the generalized Jackiw-Rossi Hamiltonian. When extended to the case of NH systems, where $\mathcal{A}$} generates a pseudo-particle-hole symmetry, the generalized NH Jackiw-Rossi operator takes the following form
@@
Notice that that the matrix operator $\Gamma_{34}$, appearing with the minimally coupled gauge field $\bm{A}$, causes the U(1) rotation between the real and imaginary components of the composite mass order, represented by $\Gamma_{3}$ and $\Gamma_{4}$, respectively. Hence, the presence of such gauge field is necessary for the stability of an isolated or deconfined vortex. The physical origin of such a gauge field and two other quantities, namely $\mu$ and $h$, depends on microscopic details. For example, in graphene-based Dirac system of spinless fermions the gauge field enters as the axial or chiral gauge field pointing in the opposite directions near two complementary valleys, and $\mu$ ($h$) corresponds to Haldane’s quantum anomalous Hall insulator mass (chiral chemical potential). On the other hand, on the surface of three-dimensional topological insulators featuring a superconducting vortex, the gauge field results from regular electromagnetic vector potential, and $\mu$ ($h$) represents the $z$ component of surface magnetization (regular chemical potential). The forthcoming discussion in this Section on the existence of zero-energy modes in the spectrum of $H^{\rm Gen}_{\rm JR}$ will be guided by the existing results in the Hermitian system ($\alpha=0$[54].
@@
The dependence of the zero modes on $\mu$ and $h$ is captured by the functions $f(r)$ and $g(r)$, which are respectively given by
@@
When $\mu>h$, as the Bessel functions are well behaved near the origin and as $r\to\infty$, the normalization of the zero mode solutions is solely determined by the exponential factors in Eq. (IV.1), thereby yielding normalizable solutions only if $|\alpha|<1$, i.e., when the effective Fermi velocity of the collection of NH Dirac fermions is real. Notice that when $h>\mu$, as the modified Bessel functions grow exponentially as $r\to\infty$, one can find normalizable solutions for zero modes if and only if $\Delta^{2}_{0}+\mu^{2}\geq h^{2}$ and when $|\alpha|<1$. The former condition is same as in Hermitian systems. Finally, I note that for $|\alpha|>1$, $J_{n}\leftrightarrow I_{n}$ and one cannot find any normalizable solution irrespective of the relative strength of $\mu$ and $h$.
@@
On the other hand, inside the vortex core the functions $f(r)$ and $g(r)$ take the form shown in Eqs. (49) and (50), respectively, with $n=1$ and $\mu^{2}-1/(2\xi^{2})\to\mu^{2}$ therein. Even in this case, I find that normalizable zero energy modes exist only when $|\alpha|<1$, i.e., when the Fermi velocity of the collection of NH Dirac fermions is real.